OpenStudy (anonymous):

In circle O, a diameteer has endpoints (-5,4) AND (3,-6). What is the length of the diameter?

OpenStudy (e.mccormick):

OK, they are trying to confuse you with all this circle talk. That diameter is just a line between two points, (-5,4) and (3,-6). So use the distance formula and that is that.\[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (anonymous):

@e.mccormick so I just plug those numbers in rgiht?

OpenStudy (anonymous):

@e.mccormick I don't have the calculator :(

OpenStudy (e.mccormick):

Yah, you plug in numbers. http://www.google.com/search?num=20&site=&source=hp&q=calculator&btnG=Search

OpenStudy (anonymous):

@e.mccormick If AB=5x-9, DC=x+3 and EF = 2x+2 what is the length of EF??

OpenStudy (e.mccormick):

Hmm... They seem to want you to solve for x and use it....

OpenStudy (e.mccormick):

Wish I had an example to work backwards from on that one. It is four unknowns in three equations, so not always solvable. If I saw a related example, I would know what they were really looking for.

OpenStudy (anonymous):